# Nets of conics containing a double line in $\mathrm{PG}(2,q)$, $q$ even

### v1

@@ Abstract
In this paper, we complete the classification of nets of conics containing at least one double line in $\mathrm{PG}(2,q)$ for $q$ even. This classification is obtained by classifying the orbits of planes in $\mathrm{PG}(5,q)$ that intersect the nucleus plane in at least one point, under the action of the group that stabilizes the Veronese surface in $\mathrm{PG}(5,q)$ for $q\geq 4$ even. As a result, we show the existence of $18$ nets containing at least one double line, $9$ of which have an empty base.
@@
In 2020, a new framework for classifying these linear systems was introduced in [18, 19], leading to significant progress. For detailed classifications and characterizations of pencils, webs, and squabs of conics, and non-empty base nets of conics in $\mathrm{PG}(2,q)$, as well as an explanation of some of the shortcomings in Campbell’s and Wilson’s partial classifications, we refer the reader to [18, 5, 19, 6, 4]. We also refer to [1], which examines nets over algebraically closed fields of characteristic zero in connection with Artinian algebras of length 7, and provides, in Appendices A and B, an interesting historical overview of nets of conics. For the classification of nets of conics over $\mathbb{R}$, we refer to the work of Wall [24].
@@
For $q$ odd, there exists a polarity of $\mathrm{PG}(5,q)$ that maps the set of conic planes of $\mathcal{V}(\mathbb{F}_{q})$ onto the set of tangent planes of $\mathcal{V}(\mathbb{F}_{q})$ (see, e.g., [15, Theorem 4.25]). This induces a correspondence between nets of conics in $\mathrm{PG}(2,q)$ containing at least one double line and planes in $\mathrm{PG}(5,q)$ that meet $\mathcal{V}(\mathbb{F}_{q})$ in at least one point, for $q$ odd. However, this correspondence does not hold over finite fields of characteristic $2$, as shown in [6]. In light of this correspondence, nets in $\mathrm{PG}(2,q)$, $q$ odd, containing at least one double line were referred to as rank-one nets. Rank-one nets of conics in $\mathrm{PG}(2,q)$, with $q$ odd, were classified in [19], resulting in $15$ orbits under $\mathrm{PGL}(3,q)$. In this paper, we use the framework from [18, 19] to classify nets of conics containing at least one double line in $\mathrm{PG}(2,q)$, $q$ even, by classifying $K$-obits of planes in $\mathrm{PG}(5,q)$ intersecting the nucleus plane non-trivially, where $K$ is the subgroup in $\mathrm{PGL}(6,q)$ stabilizing the Veronese surface $\mathcal{V}(\mathbb{F}_{q})$.
@@
There are $18$ $\mathrm{PGL}(3,q)$-orbits of nets of conics in $\mathrm{PG}(2,q)$ containing at least one double line for $q$ even.
@@ 2.2 Correspondence and combinatorial invariants
  • $\mathcal{H}_{1}$: corresponds to double lines; these hyperplanes intersect $\mathcal{V}(\mathbb{F}_{q})$ in a conic;
  • $\mathcal{H}_{2,r}$: corresponds to pairs of real lines; these hyperplanes intersect $\mathcal{V}(\mathbb{F}_{q})$ in two conics;
  • $\mathcal{H}_{2,i}$: corresponds to conjugate pairs of imaginary lines; these intersect $\mathcal{V}(\mathbb{F}_{q})$ in a single point;
  • $\mathcal{H}_{3}$: corresponds to non-singular conics; these intersect $\mathcal{V}(\mathbb{F}_{q})$ in a normal rational curve of degree $4$.
@@
Table 1: $K$-orbits of planes in $\mathrm{PG}(5,q)$ intersecting $\pi_{\mathcal{N}}$ in at least one point and their point-orbit distributions for $q>2$ even. The parameter $c$ in $\Sigma_{18}$ is a non-admissible element in $\mathbb{F}_{q}$ such that $\operatorname{Tr}(c^{-1})=\operatorname{Tr}(1)$. The parameters in $\Sigma_{20}$ satisfy $b\neq 1$ and $\operatorname{Tr}\left(c/(1+b^{2})\right)=1$. The parameter $a$ in $\Sigma_{21}$ and $\Sigma_{23}$ satisfies $\operatorname{Tr}(a)=1$.
@@
In this section, we classify the $K$-orbits of planes in $\mathrm{PG}(5,q)$ that contain no rank-$1$ points and at least one rank-$2$ point in $\pi_{\mathcal{N}}$.

### v2

@@ Abstract
This paper completes the classification of nets of conics containing at least one double line in $\mathrm{PG}(2,q)$ for $q$ even. This classification contributes to the classification of partially symmetric tensors in $\mathbb{F}_{q}^{3}\otimes S^{2}\mathbb{F}_{q}^{3}$, $q$ even. The proof is obtained using geometric and combinatorial properties of the Veronese surface in 5-dimensional projective space over the finite field of even order. In particular, the orbits of planes in $\mathrm{PG}(5,q)$ that intersect the nucleus plane of the Veronese surface in at least one point are classified. As a result, it is shown that there are exactly $18$ equivalence classes of nets in $\mathrm{PG}(2,q)$, $q$ even, containing at least one double line, $9$ of which have an empty base.
@@
In 2020, a new framework for classifying these linear systems was introduced in [18, 19], combining algebraic methods with geometric and combinatorial insights, which has since led to significant progress. For the classification and characterization of pencils, webs, and squabs of conics, and non-empty base nets of conics in $\mathrm{PG}(2,q)$, as well as an explanation of some of the shortcomings in Campbell’s and Wilson’s partial classifications, we refer the reader to [18, 5, 19, 6, 4]. We also refer to [1], which examines nets over algebraically closed fields of characteristic zero in connection with Artinian algebras of length 7, and provides, in Appendices A and B, an interesting historical overview of study of nets of conics. For the classification of nets of conics over $\mathbb{R}$, we refer to the work of Wall [24].
@@
For $q$ odd, there exists a polarity of $\mathrm{PG}(5,q)$ that maps the set of conic planes of $\mathcal{V}(\mathbb{F}_{q})$ onto the set of tangent planes of $\mathcal{V}(\mathbb{F}_{q})$ (see, e.g., [15, Theorem 4.25]). This induces a correspondence between nets of conics in $\mathrm{PG}(2,q)$ containing at least one double line and planes in $\mathrm{PG}(5,q)$ that meet $\mathcal{V}(\mathbb{F}_{q})$ in at least one point, for $q$ odd. However, this correspondence does not hold over finite fields of characteristic $2$, as shown in [6, Section 4]. In light of this correspondence, nets in $\mathrm{PG}(2,q)$, $q$ odd, containing at least one double line were referred to as rank-one nets. Rank-one nets of conics in $\mathrm{PG}(2,q)$, with $q$ odd, were classified in [19], resulting in $15$ orbits under $\mathrm{PGL}(3,q)$.
In this paper, we complete the classification of nets of conics containing at least one double line in $\mathrm{PG}(2,q)$, $q$ even, by classifying $K$-obits of planes in $\mathrm{PG}(5,q)$ intersecting the nucleus plane non-trivially, where $K$ is the subgroup in $\mathrm{PGL}(6,q)$ stabilizing the Veronese surface $\mathcal{V}(\mathbb{F}_{q})$. Moreover, for each of the orbits we determine the so-called point-orbit distributions, which are combinatorial invariants for these orbits.
@@
There are $18$ $\mathrm{PGL}(3,q)$-orbits of nets of conics in $\mathrm{PG}(2,q)$ containing at least one double line for $q$ even. Representatives and point-orbit distributions of the corresponding orbits of planes in $\mathrm{PG}(5,q)$ are as listed in Table 1.
@@ 2.2 Correspondence and combinatorial invariants
  • $\mathcal{H}_{1}$: the set of hyperplanes corresponding to double lines; these hyperplanes intersect $\mathcal{V}(\mathbb{F}_{q})$ in a conic;
  • $\mathcal{H}_{2,r}$: the set of hyperplanes corresponding to pairs of real lines; these hyperplanes intersect $\mathcal{V}(\mathbb{F}_{q})$ in two conics;
  • $\mathcal{H}_{2,i}$: the set of hyperplanes corresponding to conjugate pairs of imaginary lines; these intersect $\mathcal{V}(\mathbb{F}_{q})$ in a single point;
  • $\mathcal{H}_{3}$: the set of hyperplanes corresponding to non-singular conics; these intersect $\mathcal{V}(\mathbb{F}_{q})$ in a normal rational curve of degree $4$.
@@
Table 1: $K$-orbits of planes in $\mathrm{PG}(5,q)$ intersecting $\pi_{\mathcal{N}}$ in at least one point and their point-orbit distributions for $q>2$ even. The parameter $c$ in $\Sigma_{18}$ is a non-admissible element in $\mathbb{F}_{q}$ (see [6, Section 2.1]) such that $\operatorname{Tr}(c^{-1})=\operatorname{Tr}(1)$. The parameters in $\Sigma_{20}$ satisfy $b\neq 1$ and $\operatorname{Tr}\left(c/(1+b^{2})\right)=1$. The parameter $a$ in $\Sigma_{21}$ and $\Sigma_{23}$ satisfies $\operatorname{Tr}(a)=1$.
@@
In this section, we classify the $K$-orbits of planes in $\mathrm{PG}(5,q)$ that contain no rank-$1$ points and at least one rank-$2$ point in $\pi_{\mathcal{N}}$. We refer the reader to [2, Tables 1 and 3] for the notation, representation and properties of the $15$ $K$-orbits of lines classified in [18].